OpenStudy (rockycatsrule):

Which answer shows this equation written in standard form? 8 – 1(x – y) = –3x + 5 A. 2x + y = –3 B. 2x – y = –3 C. –4x + y = –3 D. 2x + y = 13

OpenStudy (cmtboy2):

i knoew the answer but do you

OpenStudy (lochana):

\[-x +y +3x = 5 -3\] this the last step before answer.

OpenStudy (rockycatsrule):

i am thinking it would be b

OpenStudy (cmtboy2):

no

OpenStudy (lochana):

no:)

OpenStudy (lochana):

can you show your work?

OpenStudy (cmtboy2):

h y

OpenStudy (rockycatsrule):

this one i am stuck on

OpenStudy (lochana):

it can't be -y. -1(x -y) means -x + y

OpenStudy (lochana):

negative number x negative number = positive number remember?

OpenStudy (rockycatsrule):

yes

OpenStudy (cmtboy2):

the answer is a lol

OpenStudy (rockycatsrule):

k

OpenStudy (lochana):

so you get 8 - x + y = -3x +5

OpenStudy (cmtboy2):

can i have a matle

OpenStudy (lochana):

and then move variables to the left side, others to right side.

OpenStudy (lochana):

@cmtboy2 that was not fair;)

OpenStudy (cmtboy2):

what was not

OpenStudy (lochana):

nvm. bye